Acme Company’s production budget for August is 19,300 units and includes the following component unit costs: direct materials, $9.00; direct labor, $11.80; variable overhead, $5.80. Budgeted fixed overhead is $50,000. Actual production in August was 21,870 units. Actual unit component costs incurred during August include direct materials, $10.00; direct labor, $11.20; variable overhead, $7.00. Actual fixed overhead was $53,300. The standard direct labor cost per unit consists of 0.5 hour of labor time at $23.6 per hour. During August, $244,944 of actual labor cost was incurred for 9,720 direct labor hours.

Required:
Calculate the labor rate variance and labor efficiency variance for August. (Indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable, and "None" for no effect (i.e., zero variance).)Labor rate variance Labor efficiency variance
Answer #1

>> Standard Rate = $ 23.60

>> Standard Hours = 21,870 * 0.5 = 10,935 Hours.

>> Actual Rate = ( $ 244,944 / 9720 ) = $ 25.20

>> Actual hours = 9720.

>> Labor Rate Variance = ( Standard Rate - Actual Rate ) * Actual Hours

>> Labor Rate Variance = ( $ 23.60 - $ 25.20 ) * 9,720

>> Labor Rate Variance = $ 15,552 Unfavorable.

>> Labor Efficiency Variance = ( Standard Hours - Actual Hours ) * Standard Rate

>> Labor Efficiency Variance = ( 10,935 - 9,720 ) * $ 23.6

>> Labor Efficiency Variance = $ 28,674 Favorable.